# A Computationally Efficient Approach to False Discovery Rate Control and Power Maximisation via Randomisation and Mirror Statistic

### v1

@@ 3 Simulations
We first repeat the simulations done in [Dai+22], to check whether we can achieve a similar performance to what was reported in the original paper. We then perform additional simulations under different scenarios, with the purpose of finding the limit of the approach, i.e. when the performance deteriorates too much, and to check in which situations our method can perform better than the alternatives.
The common strategy to all simulations is to perform variable selection with LASSO and inference with a standard linear model. For all simulations we set $f(a,b)=a+b$ in Eq. 2, as done in [Dai+22].
@@ 3.1 Replication of the simulations from the original paper with DS and MDS
Throughout the paper the correlation coefficient $\rho$ represents the highest value from which the correlation matrix is built, following the structure defined in Eq. 9, unless otherwise specified.
@@
Replicating the strategy used by [Dai+22] we first compare the performance of the three algorithms fixing the signal strength to $\delta=5$, varying the degree of correlation $\rho$, and then fixing the correlation to $\rho=0.5$ and varying the signal strength. MDS is run using $50$ replications of DS, as recommended by the authors.
@@
In Figures 3.1.1 and 3.1.2, top plots, we can observe some common patterns: $a)$ all three algorithms achieve, on average, FDR control at the nominal level of $0.1$; $b)$ MDS is always more conservative, achieving on average a lower FDR.
@@
Something that was missing in the original paper is a mentioning about the variability of the estimates. Here we can see that when we account for that, MDS is not much superior to the simple DS, raising some doubts about the advantage of running a much slower and computationally intensive method, compared to a simpler one.
Figure 3.1.1: Paper replication study - $\rho=0.5$
@@
Figure 3.1.2: Paper replication study - signal strength $\delta=5$
@@ 3.2.1 $\beta$ from fixed pool
Keeping all the other simulation settings as before, we run the same simulations. From Figure 3.2.1 (top) we see that the median values and variability of FDR are very similar to the ones obtained before, validating the ability of the Mirror Statistic to control the FDR. On the other hand, the results for the TPR in Figure 3.2.1 (bottom) suggest that variable selection with fixed coefficients is somewhat easier, as expected, which is reflected in higher median TPR values, concentrated near $1$.
@@
Figure 3.2.1: $\beta$ from fixed pool
@@ 3.2.2 Higher percentages of non-zero coefficients
A natural question arises regarding whether DS, MDS and RandMS, can cope with a higher percentages of active variables, potentially highly correlated. To this end, we increase the complexity of the simulated data by increasing the percentage of active variables and allowing very high degrees of correlations across the covariates.
We start by increasing the percentage of active variables to $10\%$. From Figure 3.2.2 (top) we see that the FDR is still under control and MDS is conservative, as before. In Figure 3.2.2 (bottom) we can observe that RandMS is achieving a TPR always at least comparable to DS and MDS. Common to all three methods is the sharp decrease in TPR for higher correlation structures.
Figure 3.2.2: Active coefficients $10\%$
@@
By increasing the percentage of active variables to $20\%$, we can appreciate a sharper difference, at the advantage of RandMS, both in terms of FDR and TPR. In Figure 3.2.3 (top) we see that DS and MDS start to loose control over the FDR, in particular MDS is no longer as conservative as before. RandMS is still able to achieve the required FDR control at the pre-specified level.
In Figure 3.2.3 (bottom) we see that RandMS outperform the competitors in terms of TPR, in particular for correlations up to, and including, $0.5$. On the other hand, DS is totally unable to retain enough power.
Figure 3.2.3: Active coefficients $20\%$
@@
Finally, setting the percentage of active variables to $30\%$, the difference is even more striking, again in favour of RandMS. Figures 3.2.4 (top) and 3.2.4 (bottom) show the results for FDR and TPR, respectively.
Figure 3.2.4: Active coefficients $30\%$
@@ 3.2.3 Different covariance matrix structure
In Figure 3.2.5 (top) we see that MDS is still conservative in terms of FDR, while DS and RandMS correctly control FDR at $10\%$. However, as we increase the percentage of active variables to $30\%$, Figure 3.2.6 (top), we see that MDS is not conservative anymore, while RandMS still works well.
In terms of TPR, the advantage of RandMS is more clear. Already with a percentage of active coefficients of $10\%$, Figure 3.2.5 (bottom), RandMS does a better job than DS and MDS, and, increasing that percentage to $30\%$, Figure 3.2.6 (bottom), RandMS totally outperforms the competitors.
Figure 3.2.5: Ill conditioned covariance - active coefficients $10\%$
@@
Figure 3.2.6: Ill conditioned covariance - active coefficients $30\%$
@@ 3.2.4 Increasing $p$
In Figure 3.2.7 we report the performance for $p_{1}=80$. RandMS is able to control the FDR at $0.1$ as required and is comparatively better than the other methods that show a higher variability. The TPR is much higher for RandMS up to, and including, a correlation factor of $0.7$, with TPR values close to $1$. The performance decreases for more extreme correlation levels as expected.
Figure 3.2.7: Block covariance matrix - $p=10000$ with $p_{1}=80$
@@ 3.3 Computational performance
We run a benchmark simulation of the computational requirements for the proposed Randomisation plus Mirror Statistic method. The benchmark, as well as all other analysis, has been done in Julia ([Bez+17], version $1.9.3$) on a Lenovo ThinkPad machine, with Linux OS and equipped with a 13th Gen Intel® Core™ i5-1345U × 12 CPU.
In Figure 3.3.1 we show the average time for an increasing number of variables $p$, with a fixed number of active variables $p_{1}=10$, sample size $n=300$ and a block diagonal covariance matrix (where each block is a full matrix). Both time and memory requirements scale linearly with the number of variables; as an example, with $p=10000$ the time required to estimate the full model is about $4$ seconds, while the memory required is about $0.7$ GB.
Figure 3.3.1: Computational benchmark of Randomisation with Mirror Statistic
@@
Right: Memory footprint in Gigabytes
@@ 4 Real data application to the identification of genes regulating fasting triglyceride levels
Elevated serum triglyceride (TG) levels in blood are strongly associated with increased risk of cardiovascular diseases (CVD). Serum TG levels can be reduced by a healthy diet, but there are also large inter-individual variation in fasting TG levels. An improved understanding of this variation would be beneficial for CVD prevention. In this example, we are interested in relating fasting TG levels to gene expression through a linear regression model. We have data from the screening visit of a randomized controlled dietary intervention trial, presented in detail in [Ulv+16]. In this trial, gene expression was measured in Peripheral blood mononuclear cells (PBMC). These are immune system cells and because they are circulating cells, they are exposed to nutrients, metabolites and peripheral tissues and may therefore reflect whole-body health. We include all individuals from whom we have both PBMC gene expressions and fasting TG levels, in total 251 individuals. The outcome is the log measurement of blood triglycerides, while we use measurements from $13967$ genes expression as covariates ($HumanHT-12$ Expression BeadChips. Pre-processed gene expression probe level intensity values).
@@
We proceed to analyse the data with our proposed method, i.e. outcome Randomisation plus Mirror Statistic, in order to identify which genes could potentially contribute to the differences in TG levels. We use LASSO for variable selection on the randomised outcome $U$ and a standard linear regression model for the coefficients estimation on the randomised outcome $V$ (Algorithm 1); we set the randomisation parameter $\gamma=1$ and we choose $f(a,b)=a+b$ for the test statistic in Eq. 2. The FDR target level is set to $0.1$.
@@ 5 Discussion
This extension is currently limited to Normally distributed outcomes, where randomisation takes a closed form analytical solution and the symmetry requirement of the regression coefficients to apply the Mirror Statistic is satisfied. It would be interesting to explore possible extensions to outcomes following arbitrary distributions and high-dimensional mixed models. [Lei+23] provides an extension of randomisation to distributions belonging to the exponential family, relying on the concept of conjugate distributions from Bayesian statistics. Their result could potentially be used, for example for a high-dimensional logistic regression. [Dai+23] extend the use of the Mirror Statistic to high-dimensional logistic regression, however, their method relies on the computationally expensive procedure of de-biasing the LASSO estimate, needed to satisfy the symmetry requirement to use the Mirror Statistic. Future efforts to combine these two extensions could be of practical interest. Another area of further improvement could be the adoption of different variable selection techniques in the first step of Algorithm 1. For example, substituting LASSO with a model that can better select highly correlated covariates, e.g. ElasticNet [ZH05].
Acknowledgements
The authors wish to thank Prof. Stine Ulven, Department of Nutrition, University of Oslo, for providing the Triglyceride dataset.
Code availability
The code used to perform all the simulations, the real data analysis and the computational performance benchmark is available at https://github.com/marcoelba/SelectiveInference.

### v2

@@ 3 Simulations
We first repeat the simulations done in Dai et al.19, to check whether we can achieve a similar performance to what was reported in the original paper. We then perform additional simulations under different scenarios, with the purpose of finding the limit of the approach, i.e. when the performance deteriorates too much, and to check in which situations our method can perform better than the alternatives.
The common strategy to all simulations is to perform variable selection with LASSO and inference with a standard linear model. For all simulations we set $f(a,b)=a+b$ in Equation 2.19
@@ 3.1 Replication of the simulations from the original paper with DS and MDS
  • Error variance $\sigma^{2}=1$
  • Randomization variance factor $\gamma=1$
  • Predictor in Equation 4 set to be $g(\textbf{X})=\textbf{X}\boldsymbol{\beta}$
Throughout the paper, the correlation coefficient $\rho$ represents the highest value from which the correlation matrix is built, following the structure defined in Equation 5, unless otherwise specified.
@@
Replicating the strategy used by Dai et al.19 we first compare the performance of the three algorithms fixing the signal strength to $\delta=5$, varying the degree of correlation $\rho$, and then fixing the correlation to $\rho=0.5$ and varying the signal strength. MDS is run using $50$ replications of DS, as recommended by the authors.
@@
In Figures 1 and 2, top plots, we can observe some common patterns: $a)$ all three algorithms achieve, on average, FDR control at the nominal level of $0.1$; $b)$ MDS is always more conservative, achieving on average a lower FDR.
@@
Figure 1: Paper replication study - $\rho=0.5$
@@
Figure 2: Paper replication study - signal strength $\delta=5$
@@ 3.2.1 $\beta$ from fixed pool
Keeping all the other simulation settings as before, we run the same simulations. From Figure 3 (top) we see that the median values and variability of FDR are very similar to the ones obtained before, validating the ability of the Mirror Statistic to control the FDR. On the other hand, the results for the TPR in Figure 3 (bottom) suggest that variable selection with fixed coefficients is somewhat easier, as expected, which is reflected in higher median TPR values, concentrated near $1$.
@@
Figure 3: $\beta$ from fixed pool
@@ 3.2.2 Higher percentages of non-zero coefficients
A natural question arises regarding whether DS, MDS and RandMS, can cope with a higher percentage of active variables, potentially highly correlated. To this end, we increase the complexity of the simulated data by increasing the percentage of active variables and allowing very high degrees of correlations across the covariates.
We start by increasing the percentage of active variables to $10\%$. From Figure 4 (top) we see that the FDR is still under control and MDS is conservative, as before. In Figure 4 (bottom) we can observe that RandMS is achieving a TPR always at least comparable to DS and MDS. Common to all three methods is the sharp decrease in TPR for higher correlation structures.
Figure 4: Active coefficients $10\%$
@@
By increasing the percentage of active variables to $20\%$, we can appreciate a sharper difference, at the advantage of RandMS, both in terms of FDR and TPR. In Figure 5 (top) we see that DS and MDS start to loose control over the FDR, in particular MDS is no longer as conservative as before. RandMS is still able to achieve the required FDR control at the pre-specified level.
In Figure 5 (bottom) we see that RandMS outperform the competitors in terms of TPR, in particular for correlations up to, and including, $0.5$. On the other hand, DS is totally unable to retain enough power.
Figure 5: Active coefficients $20\%$
@@
Finally, setting the percentage of active variables to $30\%$, the difference is even more striking, again in favour of RandMS. Figures 6 (top) and 6 (bottom) show the results for FDR and TPR, respectively.
Figure 6: Active coefficients $30\%$
@@ 3.2.3 Different covariance matrix structure
In Figure 7 (top) we see that MDS is still conservative in terms of FDR, while DS and RandMS correctly control FDR at $10\%$. However, as we increase the percentage of active variables to $30\%$, Figure 8 (top), we see that MDS is not conservative anymore, while RandMS still works well.
In terms of TPR, the advantage of RandMS is more clear. Already with a percentage of active coefficients of $10\%$, Figure 7 (bottom), RandMS does a better job than DS and MDS, and, increasing that percentage to $30\%$, Figure 8 (bottom), RandMS totally outperforms the competitors.
Figure 7: Ill conditioned covariance - active coefficients $10\%$
@@
Figure 8: Ill conditioned covariance - active coefficients $30\%$
@@ 3.2.4 Increasing $p$
In Figure 9 we report the performance for $p_{1}=80$. RandMS is able to control the FDR at $0.1$ as required and is comparatively better than the other methods that show a higher variability. The TPR is much higher for RandMS up to, and including, a correlation factor of $0.7$, with TPR values close to $1$. The performance decreases for more extreme correlation levels as expected.
Figure 9: Block covariance matrix - $p=10000$ with $p_{1}=80$
@@ 3.3 Computational performance
We run a benchmark simulation of the computational requirements for the proposed Randomisation plus Mirror Statistic versus Multiple Data Splitting (run with $50$ splits). The benchmark, as well as all other analysis, has been done in Julia,23, version $1.9.3$, on a Lenovo ThinkPad machine, with Linux OS and equipped with a 13th Gen Intel® Core™ i5-1345U × 12 CPU.
In Figure 10 we show the average time (in seconds) and the memory requirements (in gigabytes) to run the algorithms on a linear regression with an increasing number of variables $p$, with a fixed number of active variables $p_{1}=30$, sample size $n=300$ and a block diagonal covariance matrix. For the Randomisation method both time and memory requirements scale linearly with the number of variables; as an example, with $p=10000$ the time required to estimate the full model is about $4$ seconds, while the memory required is about $0.7$ GB. In comparison, Multiple Data Splitting requires an order of magnitude more of time and memory. For example, with $p=10000$ variables, MDS takes on average $45$ seconds and requires over $7$ GBs of memory.
Figure 10: Computational benchmark of Randomisation with Mirror Statistic vs Multiple Data Splitting
@@
Right: Memory requirement in Gigabytes
@@ 4 Real data application to the identification of genes regulating fasting triglyceride levels
Elevated serum triglyceride (TG) levels in blood are strongly associated with increased risk of cardiovascular diseases (CVD). Serum TG levels can be reduced by a healthy diet, but there are also large inter-individual variation in fasting TG levels. An improved understanding of this variation would be beneficial for CVD prevention. In this example, we are interested in relating fasting TG levels to gene expression through a linear regression model. We have data from the screening visit of a randomized controlled dietary intervention trial, presented in detail in Ulven et al.24. In this trial, gene expression was measured in Peripheral blood mononuclear cells (PBMC). These are immune system cells and because they are circulating cells, they are exposed to nutrients, metabolites and peripheral tissues and may therefore reflect whole-body health. We include all individuals from whom we have both PBMC gene expressions and fasting TG levels, in total 251 individuals. The outcome is the log measurement of blood triglycerides, while we use measurements from $13967$ genes expression as covariates ($HumanHT-12$ Expression BeadChips. Pre-processed gene expression probe level intensity values).
@@
We proceed to analyse the data with our proposed method, i.e. outcome Randomisation plus Mirror Statistic, in order to identify which genes could potentially contribute to the differences in TG levels. We use LASSO for variable selection on the randomised outcome $U$ and a standard linear regression model for the coefficients estimation on the randomised outcome $V$ (Algorithm 1); we set the randomisation parameter $\gamma=1$ and we choose $f(a,b)=a+b$ for the test statistic in Equation 2. The FDR target level is set to $0.1$.
@@ 5 Discussion
This extension is currently limited to Normally distributed outcomes, where randomisation takes a closed form analytical solution and the symmetry requirement of the regression coefficients to apply the Mirror Statistic is satisfied. It is therefore possible to use the method for the inference on linear regression models and Gaussian graphical models, for example. It would be interesting to explore possible extensions to outcomes following arbitrary distributions and high-dimensional mixed models. Leiner et al.25 provides an extension of randomisation to distributions belonging to the exponential family, relying on the concept of conjugate distributions from Bayesian statistics. Their result could potentially be used, for example for a high-dimensional logistic regression model. Dai et al.26 extend the use of the Mirror Statistic to high-dimensional logistic regression, however, their method relies on the computationally expensive procedure of de-biasing the LASSO estimate, needed to satisfy the symmetry requirement to use the Mirror Statistic. Future efforts to combine these two extensions could be of practical interest. Another area of further improvement could be the adoption of different variable selection techniques in Algorithm 1. For example, substituting LASSO with a model that can better select highly correlated covariates, e.g. ElasticNet7.